Find the equation of line passing through the point of intersection of the lines y=x and x+y=6 and which is perpendicular to 3x+6y=12
Answer by lwsshak3(11628)

Find the equation of line passing through the point of intersection of the lines y=x and x+y=6 and which is perpendicular to 3x+6y=12
**
y=x
x+y=6
..
x+x=6
2x=6
x=3
y=3
point of intersection: (3,3)
..
Given line
3x+6y=12
6y=-3x+12
y=-3x/6+12/6
y=-x/2+2
slope,m=-1/2
..
slope of line perpendicular to slope of given line=2 (negative reciprocal)
equation:
y=mx+b
y=2x+b
solve for b using coordinates of the point of intersection(3,3)
3=2*3+b
b=-3
Equation: y=2x-3 (ans.)